# Higher Steklov eigenvalues of graphs on surfaces

### v1

@@ Abstract
In this paper, we study the higher Steklov eigenvalues of graphs on surfaces. We obtain the upper bound of higher Steklov eigenvalues of a finite graph $G$ with boundary $B$ and genus $g$ by using metrical deformation via probability flows. This result can be regarded as a discrete analogue of Karpukhin’s bound in spectral geometry.
@@ 1 Introduction
These eigenvalues are called Steklov eigenvalues. Some results of domains in Euclidean spaces and Riemannian manifolds were given in [weinstock_inequalities_1954, escobar_geometry_1997, escobar_isoperimetric_1999, brock_isoperimetric_2001, wang2009sharp, karpukhin2014multiplicity, girouard2016steklov, yang2017estimates, fall2017profile, fraser2019shape, chen2024upper, girouard2024large]. Different from the Dirichlet and Neumann problems, the eigenvalue $\sigma$ for the Steklov problem is in the boundary condition.
In this paper, we study the Steklov problem on graphs, which was introduced by [hua_first_2017, hassannezhad_higher_2020]. They also introduced the Cheeger-type isoperimetric constants to estimate the first Steklov eigenvalues on graphs. He and Hua [he_upper_2022] considered the problem on trees and showed that $\sigma_{2}\leq\frac{4(D-1)}{|B|}$ and $\sigma_{2}\leq\frac{2}{L}$ where $B$, $D$ and $L$ respectively denote the boundary, maximum degree and diameter of the tree. They further gave a result on higher Steklov eigenvalues of trees. For a finite subgraph of integer lattice, Han and Hua [han_steklov_2023] proved a discrete version of Brock’s result [brock_isoperimetric_2001]. Recently, Lin and Zhao [lin_first_2024] obtained that $\sigma_{2}\leq\frac{8\Delta}{\lvert B\rvert}$ for planar graph by using circle packing and conformal mapping. They also extended He and Hua’s results to block graphs. Lin, Liu, You and Zhao  [lin2024upperboundsstekloveigenvalues] further considered $\sigma_{2}$ of graphs on surface. Some other results on upper bounds, lower bounds, Cheeger-type inequalities, rigidity problems and so on can be found in [perrin_lower_2019, perrin_isoperimetric_2021, shi_lichnerowicz-type_2022, tschanz2022upper, yu2022minimalstekloveigenvaluescombinatorial, hua2024cheegertypeinequalitiesassociated, lin2024maximizestekloveigenvaluetrees, Yu2024Monotonicity, shi2025extensionrigidityperrinslower].
The (orientable) genus $g$ of a graph $G$ is the minimum genus of an (orientable) surface in which the graph can be embedded. We generalize Lin, Liu, You and Zhao’s result as follows.
@@
Let $G=(V,E)$ be a finite graph with boundary $B$, and its maximum degree is $D$. Then the $k$-th Steklov eigenvalue
@@
We remark that the $(\mathrm{log}~g)^{2}$ factor in (1) comes from a certain geometric decomposability property of genus $g$ graphs (see Theorem 2.9) and is most likely non-essential. As a discrete analogue of Karpukhin’s bound [karpukhin2017BoundsLaplaceSteklov, Theorem 1.4] for a compact orientable Riemann surface $\Omega$ of genus $g$ with boundary $\partial\Omega$ which is shown as follows
@@
where $b$ is the number of boundary components, it is a natural problem whether $(\mathrm{log}~g)^{2}$ could be removed and $k^{2}$ could be replaced by $k$.
@@ 2.1 Steklov eigenvalues on graphs
A graph with boundary is a pair $(G,B)$, where $G=(V,E)$ is a simple connected undirected graph and $B\subset V$ is a subset of $V$ such that $\partial\left(V-B\right)=B$. Here $\partial S\coloneqq\left\{x\in V-S\mid x\sim y\ \text{for some}\ y\in S\right\}$. We call $B$ the boundary and $\Omega\coloneqq V-B$ the interior of the graph. We denote by $E(\Omega_{1},\Omega_{2})=\{\{x,y\}\in E\mid x\in\Omega_{1},y\in\Omega_{2}\}$ for $\Omega_{1},\Omega_{2}\subseteq V$. Note that the notion of graphs with boundary here is more general than the definition in [perrin_lower_2019] since the condition $E(B,B)=\emptyset$ is not required. Steklov eigenvalues of such kind of more general graphs with boundary were considered in [Colbois2014spectralgap, Colbois2018discretizations, Yu2024Monotonicity]. In this paper, we assume that $|B|\geq 2$.
@@
$(\Delta f)(x)\coloneqq\sum_{y\sim x}(f(x)-f(y)),\quad x\in V,$
@@
where $\frac{\partial f}{\partial n}(x)\coloneqq\sum\limits_{y\sim x}(f(x)-f(y))$ for $x\in B$ and $y\in V$. The reason to define $\frac{\partial f}{\partial n}$ is based on discrete Green’s formula. $\sigma\in\mathbb{R}$ is called the Steklov eigenvalue of the graph with boundary $B$. For any function $0\neq f\in\mathbb{R}^{V}$, the Rayleigh quotient of $f$ is defined as
$R(f)\coloneqq\frac{\langle f,\Delta f\rangle}{\langle f,f\rangle_{B}}=\frac{\sum\limits_{(x,y)\in E}(f(x)-f(y))^{2}}{\sum\limits_{x\in B}f^{2}(x)},$
@@
$\sigma_{k}$ $=\min_{W\subset\mathbb{R}^{V},\dim W=k}\max_{f\in W}R(f),$
where $1\leq k\leq|B|$. We can also use the following equation to characterize the Steklov eigenvalue problem on graphs:
$\Delta f=\sigma\widetilde{f}$, (2)
where $\widetilde{f}$ is defined as
$\widetilde{f}(x)\coloneqq\begin{cases}0,&x\in\Omega,\\ f(x),&x\in B.\end{cases}$
Here, we provide a lemma about the angle between subspaces and vectors first.
Lemma 2.1.
Let $v_{1},\ldots,v_{k}$ be a collection of vectors in $\mathbb{R}^{n}$ such that $v_{i}\perp v_{j}$ for any $i\neq j$ and $2\leq k\leq n$. Then for every (k1)(k-1)-dimensional subspace $W\subseteq\mathbb{R}^{n}$, there exists some $i_{0}$ such that
$\mathrm{sin}^{2}\theta_{i_{0}}\geq\frac{1}{k},$
where $\theta_{i_{0}}$ is the angle between $W$ and $v_{i_{0}}$.
Proof.
Without loss of generality, suppose that $W$ is a (k1)(k-1)-dimensional subspace defined as follows
$W\coloneqq\{(w(1),\ldots,w(n))^{\top}\in\mathbb{R}^{n}:w(1)=\ldots=w(n-k+1)=0\},$
and $v_{1},\ldots,v_{k}$ are unit vectors. Moreover, we can find some unit vectors $t_{1},\ldots,t_{n-k}$ such that $v_{1},\ldots,v_{k},t_{1},\ldots,t_{n-k}$ form a standard orthogonal basis. Let $M$ be the matrix consisting of these column vectors, that is
$M=[v_{1},\ldots,v_{k},t_{1},\ldots,t_{n-k}].$
Note that $M$ is an orthogonal matrix. Hence we have $MM^{\top}=I$, which leads to
$v_{1}(j)^{2}+\ldots+v_{k}(j)^{2}+t_{1}(j)^{2}+\ldots+t_{n-k}(j)^{2}=1,$
for any $j$. Then
$\sum_{j=1}^{n-k+1}\left(v_{1}(j)^{2}+\ldots+v_{k}(j)^{2}+t_{1}(j)^{2}+\ldots+t_{n-k}(j)^{2}\right)=n-k+1.$
Since $t_{1},\ldots,t_{n-k}$ are unit vectors, we have
$\sum_{j=1}^{n-k+1}\left(v_{1}(j)^{2}+\ldots+v_{k}(j)^{2}\right)\geq(n-k+1)-(n-k)=1.$
Therefore, there exists some $i_{0}$ such that
$\sum_{j=1}^{n-k+1}v_{i_{0}}(j)^{2}\geq\frac{1}{k}.$
The result follows immediately by $\mathrm{sin}^{2}\theta_{i_{0}}=\lVert v_{i_{0}}-P_{W}v_{i_{0}}\rVert^{2}_{2}=\sum\limits_{j=1}^{n-k+1}v_{i_{0}}(j)^{2}$, where $P_{W}$ is the projection operator onto $W$. ∎
Combining with the variational characterization of the Steklov eigenvalues, we prove the following lemma which is useful for us to get the upper bounds of higher Steklov eigenvalues.
@@
We can find Steklov eigenvectors $f_{1},\ldots,f_{|B|}$ such that
$\Delta f_{i}=\sigma_{i}\widetilde{f_{i}},\quad\forall 1\leq i\leq|B|,$
and $f_{1}|_{B},\ldots,f_{|B|}|_{B}$ is an orthogonal basis.
By the definition, $\mathrm{span}\{v_{1},\ldots,v_{k}\}$ is a $k$-dimensional subspace in $\mathbb{R}^{V}$ and $\mathrm{span}\{v_{1}|_{B},\ldots,v_{k}|_{B}\}$ is a $k$-dimensional subspace in $\mathbb{R}^{B}$. By Lemma 2.1, there exists some $i_{0}$ such that
$\mathrm{cos}^{2}(\frac{\pi}{2}-\theta)=\mathrm{sin}^{2}\theta\geq\frac{1}{k},$
where $\theta$ is the angle between $v_{i_{0}}|_{B}$ and the subspace $\mathrm{span}\{f_{1}|_{B},\ldots,f_{k-1}|_{B}\}$. Note that $\pi/2-\theta$ is the angle between $v_{i_{0}}|_{B}$ and the subspace $\mathrm{span}\{f_{k}|_{B},\ldots,f_{|B|}|_{B}\}$. We write $v_{i_{0}}|_{B}$ as
$v_{i_{0}}|_{B}=\sum_{1\leq j\leq k-1}c_{j}f_{j}|_{B}+\sum_{k\leq l\leq|B|}c_{l}f_{l}|_{B}.$
Then there exists $\delta\in\mathbb{R}^{V}$ such that $\delta|_{B}=0$ and
$v_{i_{0}}=\sum_{1\leq j\leq k-1}c_{j}f_{j}+\sum_{k\leq l\leq|B|}c_{l}f_{l}+\delta.$
Since $f_{j}$ ($1\leq j\leq|B|$) and $\delta$ are pairwise orthogonal on $B$ and $\Delta$ is a self-adjoint positive semi-definite operator, we have
$R(v_{i_{0}})=$ $\frac{\langle v_{i_{0}},\Delta v_{i_{0}}\rangle}{\langle v_{i_{0}},v_{i_{0}}\rangle_{B}}=\frac{\langle v_{i_{0}},\Delta v_{i_{0}}\rangle}{\langle v_{i_{0}}|_{B},v_{i_{0}}|_{B}\rangle}$
$=$ $\frac{1}{\lVert v_{i_{0}}|_{B}\rVert_{2}^{2}}\left(\sum_{1\leq j\leq k-1}\langle c_{j}f_{j}|_{B},\sigma_{j}c_{j}f_{j}|_{B}\rangle+\sum_{k\leq l\leq|B|}\langle c_{l}f_{l}|_{B},\sigma_{l}c_{l}f_{l}|_{B}\rangle+\langle\delta,\Delta\delta\rangle\right)$
$\geq$ $\frac{1}{\lVert v_{i_{0}}|_{B}\rVert_{2}^{2}}\sum_{k\leq l\leq|B|}\langle c_{l}f_{l}|_{B},\sigma_{l}c_{l}f_{l}|_{B}\rangle$
$=$ $\sum_{k\leq l\leq|B|}\sigma_{l}\frac{\lVert c_{l}f_{l}|_{B}\rVert_{2}^{2}}{\lVert v_{i_{0}}|_{B}\rVert_{2}^{2}}$
$\geq$ $\sigma_{k}\frac{\sum_{k\leq l\leq|B|}\lVert c_{l}f_{l}|_{B}\rVert_{2}^{2}}{\lVert v_{i_{0}}|_{B}\rVert_{2}^{2}}$
$=$ $\sigma_{k}\mathrm{cos}^{2}(\frac{\pi}{2}-\theta)$
$\geq$ $\frac{\sigma_{k}}{k}.$
Therefore, $\frac{\sigma_{k}}{k}\leq\max\limits_{1\leq i\leq k}\{R(v_{i})\}$. ∎
@@ 2.3 Padded decomposition
We define the diameter of $G$ in $(V,d_{\omega})$ as $diam(G)\coloneqq\max\limits_{u,v\in V(G)}d_{\omega}(u,v)$.
Let $(X,d)$ be a finite semi-metric space. We recall the definitions of the padded decomposition and the modulus of padded decomposition.
@@
Let $(X,d)$ be a finite semi-metric space. If $P$ is a partition of $X$, we will also regard it as a function $P:X\rightarrow 2^{X}$ such that for every $x\in X$, $P(x)$ is the unique part $C\in P$ for which $x\in C$. Let $\nu$ be a probability distribution over partitions of $X$, and $P$ a random partition distributed according to $\nu$. We say that $P$ is $\kappa$-bounded if it holds that for every $C\in P$, we have $diam(C)\leq\kappa$. We say that $P$ is $(\alpha,\delta)$-padded if
@@
The modulus $\alpha(X,d;\delta,\kappa)$ of padded decomposibility of a finite semi-metric space $(X,d)$ is defined as
@@
Now we state two important theorems about the modulus of padded decomposibility in [biswal2010eigenvalue] and [lee2010genus].
Theorem 2.8 ([biswal2010eigenvalue, Corollary 4.2]).
@@ 3.1 Spreading weights
$\sigma_{k}\leq\frac{64k}{\delta}\frac{D}{|B|}(\frac{\alpha}{\epsilon})^{2},$
@@
$\sum_{uv\in E}|f_{i}(u)-f_{i}(v)|^{2}=$ $\sum_{u\in\tilde{S}_{i}}\sum_{v:uv\in E}|f_{i}(u)-f_{i}(v)|^{2}$
$\leq$ $\sum_{u\in\tilde{S}_{i}}\sum_{v:uv\in E}d_{\omega}(u,v)^{2}$
$=$ $\sum_{u\in\tilde{S}_{i}}\sum_{v:uv\in E}\left[\omega(u)+\omega(v)\right]^{2}$
@@
Before our proof, we recall Slater’s condition in convex optimization, which can be found in Chapter 5 of [boyd2004convex].
@@
According to the theory of Lagrange dual shown in [boyd2004convex], $(\bm{P})$ and its dual $(\bm{P^{*}})$ have the following equivalent forms:
@@
$\sigma_{k}\leq\frac{64kD}{\delta|B|}(\frac{\alpha}{\epsilon})^{2},$
@@
$\sigma_{k}\lesssim\frac{kD(\mathrm{log}~g)^{2}}{(\epsilon_{\lfloor|B|/32k\rfloor}(G,B,\omega))^{2}|B|}$ (5)
@@
$\sigma_{k}\lesssim\frac{kD(\mathrm{log}~g)^{2}}{(\epsilon_{r}(G,B,\omega_{r}))^{2}|B|}\lesssim\frac{kDg(\mathrm{log}~g)^{2}}{r}\lesssim Dg(\mathrm{log}~g)^{2}\frac{k^{2}}{|B|},$
for $r=\lfloor|B|/32k\rfloor\geq C_{0}\sqrt{g}$. If $\lfloor|B|/32k\rfloor<C_{0}\sqrt{g}$ or $k\geq\delta\lvert B\rvert/4$, then $\sigma_{k}=\mathcal{O}(Dg(\mathrm{log}~g)^{2}\frac{k^{2}}{|B|})$ holds trivially due to the bound $\sigma_{k}\leq kD$ obtained by Lemma 2.2 for all $1\leq k\leq|B|$.
Above all, our proof is complete. ∎

### v2

@@ Abstract
In this paper, we study the higher Steklov eigenvalues of graphs on surfaces. We obtain the upper bound of higher Steklov eigenvalues of a finite graph $G$ with boundary $B$ and genus $g$ by using metrical deformation via probability flows. Our result can be regarded as a discrete analogue of Karpukhin’s bound in spectral geometry. Moreover, this result implies the upper bound of higher Laplacian eigenvalues given by Kelner, Lee, Price and Teng (Geom. Funct. Anal., 2011).
@@ 1 Introduction
These eigenvalues are called Steklov eigenvalues. Some results of domains in Euclidean spaces and Riemannian manifolds were given in [31, 7, 8, 3, 30, 18, 11, 32, 9, 10, 4, 12]. Different from the Dirichlet and Neumann problems, the eigenvalue $\sigma$ for the Steklov problem is in the boundary condition.
In this paper, we study the Steklov problem on graphs, which was introduced by [16, 14]. The idea of obtaining the upper bound for the Escobar Cheeger-type constant introduced in [16] motivates us to find the upper bounds for the Steklov eigenvalues of graphs. He and Hua [15] considered the problem on trees and showed that $\sigma_{2}\leq\frac{4(D-1)}{|B|}$ and $\sigma_{2}\leq\frac{2}{L}$, where $B$, $D$, and $L$ respectively denote the boundary, maximum degree, and diameter of the tree. They further provided a result on higher Steklov eigenvalues of trees. For a finite subgraph of the integer lattice, Han and Hua [13] proved a discrete version of Brock’s result [3]. Recently, Lin and Zhao [24] obtained that $\sigma_{2}\leq\frac{8D}{\lvert B\rvert}$ for planar graphs by using circle packing and conformal mapping. They also extended He and Hua’s results to block graphs. Lin, Liu, You, and Zhao  [22] further considered $\sigma_{2}$ of graphs on surfaces. Some other results on upper bounds, lower bounds, Cheeger-type inequalities, rigidity problems, and so on can be found in [25, 26, 27, 29, 33, 17, 23, 34, 28].
The (orientable) genus $g$ of a graph $G$ is the minimum genus of an (orientable) surface in which the graph can be embedded. In this paper, we generalize Lin, Liu, You and Zhao’s result as follows.
@@
Let $G=(V,E)$ be a finite connected graph with boundary $B$, and its maximum degree is $D$. Then the $k$-th Steklov eigenvalue
@@
Our result also implies Kelner, Lee, Price and Teng’s result [20] if we choose $B=V$. We remark that the $(\mathrm{log}~g)^{2}$ factor in (1) comes from a certain geometric decomposability property of genus $g$ graphs (see Theorem 2.8) and is most likely non-essential. As a discrete analogue of Karpukhin’s bound [19, Theorem 1.4] for a compact orientable Riemann surface $\Omega$ of genus $g$ with boundary $\partial\Omega$ which is shown as
@@
where $b$ is the number of boundary components, it is a natural problem whether $(\mathrm{log}~g)^{2}$ could be removed and $gk$ could be replaced by $g+k$.
@@ 2.1 Steklov eigenvalues on graphs
Let $G=(V,E)$ be an undirected simple graph, where $V,E$ are respectively the set of vertices and edges of $G$. A graph with boundary is a pair $(G,B)$, where $B\subset V$ is a non-empty set. The set $B$ is called the boundary and $\Omega\coloneqq V\setminus B$ is called the interior. We denote by $E(\Omega_{1},\Omega_{2})=\{\{x,y\}\in E\mid x\in\Omega_{1},y\in\Omega_{2}\}$ for $\Omega_{1},\Omega_{2}\subseteq V$. Notice that the notion of graphs with boundary here is more general than the definition in [25] since the condition $E(B,B)=\emptyset$ is not required and the boundary vertices can be not adjacent to any interior vertices. Steklov eigenvalues of such more general graphs with boundary were considered in [6, 5, 34]. In this paper, we assume that $|B|\geq 2$ and $G$ is connected.
@@
$(\Delta f)(x)\coloneqq\sum_{\{x,y\}\in E}(f(x)-f(y)),\quad x\in V,$
@@
where $\frac{\partial f}{\partial n}(x)\coloneqq\sum\limits_{\{x,y\}\in E}(f(x)-f(y))=-\Delta f(x)$ for $x\in B$ and $y\in V$. The reason to define $\frac{\partial f}{\partial n}$ is based on discrete Green’s formula. When $B=V$, $\sigma$ is the Laplacian eigenvalue of $G$. Here $\sigma\in\mathbb{R}$ is called the Steklov eigenvalue of the graph with boundary $B$. For any function $0\neq f\in\mathbb{R}^{V}$, the Rayleigh quotient of $f$ is defined as
$R_{(G,B)}(f)\coloneqq\frac{\langle f,\Delta f\rangle}{\langle f,f\rangle_{B}}=\frac{\sum\limits_{\{x,y\}\in E}(f(x)-f(y))^{2}}{\sum\limits_{x\in B}f^{2}(x)},$
@@
$\sigma_{k}(G,B)$ $=\min_{W\subset\mathbb{R}^{V},\dim W=k}\max_{f\in W}R_{(G,B)},$
where $1\leq k\leq|B|$.
The following lemma is useful to obtain the upper bounds of higher Steklov eigenvalues.
@@
By the definition, $W=\mathrm{span}\{v_{1},\ldots,v_{k}\}$ is a $k$-dimensional subspace. For any $f=\sum_{i=1}^{k}c_{i}v_{i}\in W$ and $x,y\in V$, we claim that
$(f(x)-f(y))^{2}\leq 2\sum_{i=1}^{k}c_{i}^{2}(v_{i}(x)-v_{i}(y))^{2}.$ (2)
If $f(x)=0$ or $f(y)=0$, there is nothing to prove. Then since $v_{i}$’s have disjoint supports, we can suppose that $p$ (resp. $q$) is the only index satisfying that $v_{p}(x)\neq 0$ (resp. $v_{q}(x)\neq 0$). In the case $p\neq q$, we have
$(f(x)-f(y))^{2}=$ $(c_{p}v_{p}(x)-c_{q}v_{q}(y))^{2}$
$\leq$ $2(c_{p}^{2}v_{p}(x)^{2}+c_{q}^{2}v_{q}(x)^{2})$
$\leq$ $2\sum_{i=1}^{k}c_{i}^{2}(v_{i}(x)-v_{i}(y))^{2}.$
The case $p=q$ is direct. Hence by (2), we get
$R_{(G,B)}(f)=$ $\frac{\sum\limits_{\{x,y\}\in E}(f(x)-f(y))^{2}}{\sum\limits_{x\in B}f^{2}(x)}$
$\leq$ $\frac{2\sum\limits_{i=i}^{k}c_{i}^{2}\sum\limits_{\{x,y\}\in E}(v_{i}(x)-v_{i}(y))^{2}}{\sum\limits_{i=i}^{k}c_{i}^{2}\sum\limits_{x\in B}v_{i}^{2}(x)}$
$\leq$ $2\max\limits_{1\leq i\leq k}\{R_{(G,B)}(v_{i})\}.$
Therefore, $\sigma_{k}(G,B)\leq 2\max\limits_{1\leq i\leq k}\{R_{(G,B)}(v_{i})\}$. ∎
@@ 2.3 Padded decomposition
We also define the diameter of $G$ in $(V,d_{\omega})$ as $diam(G)\coloneqq\max\limits_{u,v\in V(G)}d_{\omega}(u,v)$. Next, we recall the definitions of the padded decomposition and the modulus of padded decomposition.
@@
Let $(X,d)$ be a finite metric space. If $P$ is a partition of $X$, we will also regard it as a function $P:X\rightarrow 2^{X}$ such that for every $x\in X$, $P(x)$ is the unique part $C\in P$ for which $x\in C$. Let $\nu$ be a probability distribution over partitions of $X$, and $P$ a random partition distributed according to $\nu$. We say that $P$ is $\kappa$-bounded if it holds that for every $C\in P$, we have $diam(C)\leq\kappa$. We say that $P$ is $(\alpha,\delta)$-padded if
@@
The modulus $\alpha(X,d;\delta,\kappa)$ of padded decomposibility of a finite metric space $(X,d)$ is defined as
@@
Now we state two important theorems about the modulus of padded decomposibility in [1] and [21].
Theorem 2.7 ([1, Corollary 4.2]).
@@ 3.1 Spreading weights
$\sigma_{k}\leq\frac{128}{\delta}\frac{D}{|B|}(\frac{\alpha}{\epsilon})^{2},$
@@
$\sum_{uv\in E}|f_{i}(u)-f_{i}(v)|^{2}=$ $\sum_{u\in\tilde{S}_{i}}\sum_{v:\{u,v\}\in E}|f_{i}(u)-f_{i}(v)|^{2}$
$\leq$ $\sum_{u\in\tilde{S}_{i}}\sum_{v:\{u,v\}\in E}d_{\omega}(u,v)^{2}$
$=$ $\sum_{u\in\tilde{S}_{i}}\sum_{v:\{u,v\}\in E}\left[\omega(u)+\omega(v)\right]^{2}$
@@
Before our proof, we recall Slater’s condition in convex optimization, which can be found in Chapter 5 of [2].
@@
According to the theory of Lagrange dual shown in [2], $(\bm{P})$ and its dual $(\bm{P^{*}})$ have the following equivalent forms:
@@
$\sigma_{k}\leq\frac{128D}{\delta|B|}(\frac{\alpha}{\epsilon})^{2},$
@@
$\sigma_{k}\lesssim\frac{D(\mathrm{log}~g)^{2}}{(\epsilon_{\lfloor|B|/32k\rfloor}(G,B,\omega))^{2}|B|}$ (5)
@@
$\sigma_{k}\lesssim\frac{D(\mathrm{log}~g)^{2}}{(\epsilon_{r}(G,B,\omega_{r}))^{2}|B|}\lesssim\frac{Dg(\mathrm{log}~g)^{2}}{r}\lesssim Dg(\mathrm{log}~g)^{2}\frac{k}{|B|},$
for $r=\lfloor|B|/32k\rfloor\geq C_{0}\sqrt{g}$. If $\lfloor|B|/32k\rfloor<C_{0}\sqrt{g}$ or $k\geq\delta\lvert B\rvert/4$, then $\sigma_{k}=\mathcal{O}(Dg(\mathrm{log}~g)^{2}\frac{k}{|B|})$ holds trivially due to the bound $\sigma_{k}\leq 2D$ obtained by Lemma 2.1 for all $1\leq k\leq|B|$.
Our proof is complete. ∎
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
No data was used for the research described in the article.